OpenStudy (anonymous):

If a person has ATM fees each month of $18 for 8 years, what would be the total cost of those bank fees

OpenStudy (anonymous):

The simple answer to your question is: ($18.00*12)*8=$1,728.00. But lets have some fun with the funds instead of giving it to the bank for the privilege of withdrawing your own money. You could create an annuity with the $18.00/month by depositing your money every month into a savings account. With the crappy savings account interest rates currently, you would typically gain interest at .05% annually. Using the Future value of an annuity calculation: FVAn= PMT*((1+i)^n-1)/i. Where PMT=monthly payment of $18.00, i=annual interest rate divided by the number of months in the year(.0005/12=.00004167), n= number of payment periods in annuity(12months*8years=96)= $1731.42 is the total for the eight years. This includes the principal payment of $1728.00 plus $3.00 in interest! I know this does not sound like a lot of money but this is also a rough time in our economy for simple risk free investing.